In April 1974, Steve Prefontaine completed a 10 km race in a time of 27 min, 43.6 s. Suppose Pre was at the 7.71 km mark at a time of 25.0 min. If he accelerated for 60 s and then maintained his increased speed for the remainder of the race, มั่ calculate his acceleration over the 60 s interval. Assume his instantaneous speed at the 7.71 km mark was the same as his overall average speed up to that time acceleration: m/s2
Answer #1

Lots of numbers, but shouldn't be too tough.
At to = 25 min, v = Vavg = 7710m / (25*60s) = 5.14 m/s
He had to cover (10 - 7.71)km = 2.29 km = 2290 m
in t = 27m43.6s - 25min = 2m43.6s = 163.6 s

After he accelerates for 60 s, he will have velocity V,
which he will run for the remaining (163.6 - 60)s = 103.6 s.
Furthermore, we know that V = v + at = 5.14m/s + a*60s

Distance covered while accelerating is
s = vt + ½at² = 5.14m/s * 60s + ½a(60s)² = 308.4m + a*1800s²
Distance covered while "coasting" is
S = Vt = (5.14m/s + a*60s)*103.6s = 532.504m + a*6216s²
We know that s + S = 1730 m, so
2290 m = 840.904m + a*8016s²
a = 1449.096m / 8016s²

a = 0.1807 m/s²